OpenStudy (anonymous):

solve the equation (3tan^2x-1)(tan^2x-3)=0

OpenStudy (anonymous):

Solve these equations: (3tan^2x-1)=0 (tan^2x-3)=0

OpenStudy (anonymous):

u kno wen tan is zero?

OpenStudy (anonymous):

tanx is zero if x = 0

OpenStudy (anonymous):

there are infinte solutions for x. What is the range of x